Can someone help me with this problem?
I need to find the prime factorization of this number:
90
Thanks,
sher

Answer by BGluvsmath(64)

Solved by pluggable solver: Factor any number
90 is NOT a prime number: 90 = 2 * 3 * 3 * 5

Work Shown

90 is divisible by 2: 90 = 45 * 2.
45 is divisible by 3: 45 = 15 * 3.
15 is divisible by 3: 15 = 5 * 3.
5 is not divisible by anything.

This calculation used this Prime Factorization Algorithm.